solve by system of addition
2x = 2 - y
3x + y = -1

Answer by checkley77(12844)

2x=2-y
or 2x+y=2 multiply by -1 & add.
3x+y=-1
-2x-y=-2
-------------------
x=-3 ans.
2*-3=2-y
-6=2=-y
-y=-6-2
-y=-8
y=8 ans.
Proof:
3*-3+8=-1
-9+8=-1
-1=-1